# Aspects of the commuting graph [^1] To the memory of Richard Parker

### v3

@@ 8.1.1 Classical Lie type
Proof 9
These finite simple groups are categorized as follows:
@@
Next, consider $\mathop{\mathrm{PSp}}\nolimits_{2m}(q)$. Recall that the symplectic group $\mathop{\mathrm{Sp}}\nolimits_{2m}(q)$ is the subgroup of $\mathop{\mathrm{GL}}\nolimits_{2m}(q)$ consisting of all isometries of a nonsingular alternating bilinear form on the vector space $\mathbb{F}_{q}^{2m}$. It turns out that $|\mathop{\mathrm{Sp}}\nolimits_{2m}(q)|>q^{m^{2}}$. As $|\mathop{\mathrm{GL}}\nolimits_{2m}(q)|=q^{4m^{2}}$, clearly
@@
Thus, the description of the group $\mathop{\mathrm{Sp}}\nolimits_{2m}(q)$ is easy to construct. Given $2m$ and $q$ in binary, in time polynomial in $m$ and $q$ we can define a nonsingular alternating bilinear form $\langle$,$\rangle$ , ⟩ on $\mathbb{F}_{q}^{2m}$. Then, as done for the previous case, we can enumerate the elements of $\mathop{\mathrm{GL}}\nolimits_{2m}(q)$, one by one, and put each into $\mathop{\mathrm{Sp}}\nolimits_{2m}(q)$ if it is an isometry for $\langle$,$\rangle$ , ⟩. Now, the simple group $\mathop{\mathrm{PSp}}\nolimits_{2m}(q)$ by quotienting out the subgroup of scalar matrices (which has size at most $2$). Clearly, Lemma 8.2 can be applied to obtain the desired description for $\mathop{\mathrm{PSp}}\nolimits_{2m}(q)$ along with the algorithm for obtaining its multiplication table.
For the next case $\mathop{\mathrm{PSU}}\nolimits_{n}(q)$, we start with the group $\mathop{\mathrm{GU}}\nolimits_{n}(q)$ which is the subgroup of $\mathop{\mathrm{GL}}\nolimits_{n}(q^{2})$ that preserves the sesquilinear form $\langle x,y\rangle=\sum_{i=1}^{n}x_{i}y_{i}^{q}$. From there on the process of obtaining $\mathop{\mathrm{SU}}\nolimits_{n}(q)$ and then $\mathop{\mathrm{PSU}}\nolimits_{n}(q)$ is exactly like in the case of $\mathop{\mathrm{PSL}}\nolimits_{n}(q)$. In order to obtain the desired description for $\mathop{\mathrm{PSU}}\nolimits_{n}(q)$ it suffices to generate $\mathop{\mathrm{GU}}\nolimits_{n}(q)$ in time polynomial in $|\mathop{\mathrm{PSU}}\nolimits_{n}(q)|$ given $n$ and $q$. Clearly, $\mathop{\mathrm{GU}}\nolimits_{n}(q)$ can be generated in time polynomial in $|\mathop{\mathrm{GL}}\nolimits_{n}(q^{2})|$ by looking all these matrices. However, $|\mathop{\mathrm{GL}}\nolimits_{n}(q^{2})|<|\mathop{\mathrm{GU}}\nolimits_{n}(q)|^{3}$. Thus, the naive algorithm for generating all of $\mathop{\mathrm{GL}}\nolimits_{n}(q^{2})$ and picking those which preserve the sesquilinear form will again work. Finally, the group $\mathop{\mathrm{PSU}}\nolimits_{n}(q)$ is only smaller than $\mathop{\mathrm{GU}}\nolimits_{n}(q)$ by a small power of $q$.
@@ 8.1.2 Exceptional Lie type
We will closely follow Wilson’s exposition [77, Chapter 4]. These finite simple groups come in ten families:
@@
    $\mathop{\mathrm{Sz}}\nolimits(2^{2n+1})$:  Let $q=2^{2n+1}$. The Suzuki group $Sz(q)$ is a subgroup of the symplectic group $\mathop{\mathrm{Sp}}\nolimits_{4}(2^{2n+1})$. It arises as the centralizer of a particular automorphism of order $2$ in $\mathop{\mathrm{Sp}}\nolimits_{4}(2^{2n+1})$. As $|\mathop{\mathrm{Sz}}\nolimits(q)|=(q^{2}+1)q^{2}(q-1)$ $|\mathop{\mathrm{GL}}\nolimits_{4}(q)|\leq|\mathop{\mathrm{Sz}}\nolimits(q)|^{4}$. Thus, we can cycle through all elements of $\mathop{\mathrm{Sp}}\nolimits_{4}(q)$ looking for the elements of this centralizer. The automorphism is also easily described (in size independent of $q$).
@@
    $G_{2}(q)$:  It turns out that $G_{2}(q)$ arises as the automorphism group of the octonion algebra over $\mathbb{F}_{q}$. The entire octonion algebra can be listed in time $O(q^{8})$. As the automorphisms permute fix the identity in the octonion algebra, $G_{2}(q)$ is essentially a subgroup of $\mathop{\mathrm{GO}}\nolimits_{7}(q)$. We can enumerate $\mathop{\mathrm{GO}}\nolimits_{7}(q)$ and pick out the automorphisms which will give us $G_{2}(q)$. As $|G_{2}(q)|$ is $q^{6}(q^{6}-1)(q^{2}-1)$, this will be polynomial time in the group size. We can define the $\mathop{\mathrm{desc}}(G_{2}(q))$ of size $O(\log q)$. It is essentially just the encoding of $q$.
@@
    $R_{1}(3^{2n+1})={}^{2}G_{2}(3^{2n+1})$: This can be handled quite analogously to $G_{2}(q)$ [77, Section 4.5].
@@
    ${}^{3}{D}_{4}(q)$:  This group arises as centralizer of a certain order $3$ automorphism of $\mathop{\mathrm{P}\Omega}\nolimits^{+}_{8}(q^{3})$. Because of the special nature of this automorphism (product of an $\mathbb{F}_{q}$ automorphism and the triality automorphism), it has a short description of size $O(\log q)$. As $\mathop{\mathrm{P}\Omega}\nolimits^{+}_{8}(q^{3})$ can be enumerated in time polynomial in $q$, we can pick out the centralizer elements in time polynomial in $q$.
@@
    $F_{4}(q)$:  when $q$ is not a power of $2$ or $3$, $F_{4}(q)$ arises as automorphisms of the finite (non-associative) Albert algebra which is $27$-dimensional. Thus, $F_{4}(q)$ can be picked out from $GL_{27}(q)$ by enumerating it and checking each element for being an automorphism. The finite Albert algebra can be defined in $O(\log q)$ size by defining the multiplication of its basis elements.
@@
    $R_{2}(2^{2n+1})={}^{2}F_{4}(2^{2n+1})$:  This case is quite similar to the previous one [77, Section 4.9].
@@
    $E_{6}(q)$:  This group can be defined as automorphisms of a certain trilinear form over the Albert algebra. After quotienting out the scalar subgroup from this automorphism group we will obtain $E_{6}(q)$. The same algorithm of enumerating the bigger group and selecting the automorphisms and then quotienting out the scalars will work.
@@
    $E_{7}(q)$ and $E_{8}(q)$:  Both these arise as automorphisms of a finite Lie algebra. For $E_{8}(q)$ we have to consider a $248$ dimensional Lie algebra. Thus we can look for $E_{8}(q)$ inside $GL_{248}(q)$ by enumerating all its elements and picking out the automorphisms. All we need is that from a $O(\log q)$ size description we should be able to obtain the Lie algebra. For that we only need to express the multiplication of basis elements as an $\mathbb{F}_{q}$ linear combination of the basis (which is of size $248$). $E_{7}(q)$ is similarly dealt with as automorphisms of a smaller dimensional Lie algebra.$\qed$
@@ 8.2 Short descriptions for finite groups
There is a polynomial-time algorithm that takes a group $G$ by its multiplication table as input and outputs a cube-generating sequence of size $2\log|G|$ for $G$.
Proof 11
@@
$C_{i}=\{g_{1}^{\epsilon_{1}}g_{2}^{\epsilon_{2}}\cdots g_{i}^{\epsilon_{k}}\mid\epsilon_{j}\in\{0,1\}\hbox{ for each }j\}.$
For any $g\in G$ let $C_{i}g=\{xg\mid x\in C_{i}\}$. We can enumerate all the elements of $C_{i}$ and $C_{i}g$ with at most $2i\cdot 2^{i}$ group multiplications, for $i\leq\log|G|$, which is is polynomial in $|G|$.
Now, if $C_{i}g\cap C_{i}\neq\emptyset$ for $g\in G$ then $g$ is in the cube generated by the sequence $(g_{1}^{-1},g_{2}^{-1},\ldots,g_{i}^{-1},g_{1},\ldots,g_{i})$. Thus, if for all $g\in G$ $C_{i}g\cap C_{i}\neq\emptyset$ then we have a cube generating sequence of length $2i$. Otherwise, there is a $g_{i+1}\in G$ such that
@@
and the cube generated by $(g_{1},g_{2},\ldots,g_{i+1})$ is twice the size of $C_{i}$ by a simple counting argument. Clearly, we can find $g_{i+1}$ in time polynomial in $|G|$ by searching through $G$$\qed$
@@
where $w[x,y]$ is a word over the cube generating sequence $Y_{i-1}$. Thus, $w(x,y)$ is a word of length at most $2\log|N_{i-1}|$, expressed using the c.g.s. $Y_{i-1}$. Let $R_{i}$ consist of the set of all such equations defined by Equation (6) for $y\in Y_{i-1}$ and $x\in X_{i}$. Notice that $R_{i}$ has $|Y_{i-1}|\cdot|X_{i}|$ equations in it and each word $w(x,y)$ over $Y_{i-1}$ is defined by the bit vector of exponents in the cube product over $Y_{i-1}$. Thus, the size of $R_{i}$ is
@@
Using the above claim repeatedly we can express products of the form $xa$ for $x\in C,a\in N_{i-1}$ as $a^{\prime}x$ for som $a^{\prime}\in N_{i-1}$. Finally, we can express the product of any two elements $a_{1}b_{1},a_{2}b_{2}\in N_{i}$, where $a_{1},a_{2}\in N_{i-1}$ and $b_{1},b_{2}\in C$ in the form $a_{3}b_{3},a_{3}\in N_{i-1},b_{3}\in C$. The algorithm runs in time polynomial in $|N_{i-1}|$ and $|X_{i}|$ and hence in time polynomial in $|N_{i}|$.
@@
    More precisely, for each $g\in G$ we compute its normal closure by starting with the cyclic subgroup $N=\langle g\rangle$ and repeatedly replacing it with the (larger) subgroup generated by the set $\{xyx^{-1},x\in G,y\in N\}$ until $N$ does not grow any further. By construction, the final $N$ is a normal subgroup of $G$ and its computation takes time polynomial in $|G|$. If $G$ is not simple, then for some $g\in G$ this construction will yield a nontrivial normal subgroup $N$. We can now recursively find composition series for $N$ and $G/N$ (we have multiplication tables for both) and put them together to obtain a composition series for $G$ in time polynomial in $|G|$.
@@
This algorithm is broadly based on a result of McIver and Neumann [56] that bounds the number of $n$ element groups by $2^{O(\log^{3}n)}$. The theorem is sharpened by subsequent work of Babai et al. [5] on short presentations for finite groups. They show that groups of order $n$ (that do not have a specific finite simple group type, called the Ree groups, as section) have short generator-relator presentations $\langle Y|R\rangle$ of size $O(\log^{3}n)$. The crux of their proof is that all simple groups of $n$ elements (except the Ree groups) have generator-relator presentations of size $O(\log^{2}n)$. Combined with the fact that $n$ element groups have composition series of $\log n$ length gives the $O(\log^{3}n)$ bound.
@@ 9 Isoclinism of groups and commuting graphs
A property of extraspecial groups that we exploited in the algorithm is that extraspecial groups of the same order are isoclinic. We saw that isoclinic groups of the same order have isomorphic commuting graphs.
What about the converse? That is, if two groups have isomorphic commuting graphs, must they be isoclinic? This holds for various classes of groups, such as abelian groups, nonabelian simple groups, and extraspecial groups (as we have seen).
However, it is not true in general; we recycle an example taken from [23] to show this. Let $G$ be the group of order $64$ which is SmallGroup(64,182) in the SmallGroups library in GAP [36]. The Schur multiplier of $G$ has order $2$, so a Schur cover (a group $H$ of maximal order subject to having a subgroup $Z\leq H^{\prime}\cap Z(H)$ such that $H/Z\cong G$) has order $128$. Moreover, the Bogomolov multiplier of $G$ is equal to the Schur multiplier, which implies that the Schur covers $H$ are commutation-preserving: that is, two elements $a,b\in H$ commute if and only if their projections $Za,Zb\in G$ commute. This implies that the commuting graph of a Schur cover is obtained from the commuting graph of $G$ by replacing each vertex with a clique of size $2$, with all edges between cliques corresponding to adjacent vertices. This procedure also describes the commuting graph of $G\times C_{2}$. On the other hand, it is easy to verify computationally that the derived groups of $H$ and $G\times C_{2}$ are not isomorphic, so these groups cannot be isoclinic.

### v4

@@ 8.1.1 Classical Lie type
Proof 17
These finite simple groups are categorized below, following Wilson’s text [79, Chapter 3].
@@
Next, consider $\mathop{\mathrm{PSp}}\nolimits_{2m}(q)$. Recall that the symplectic group $\mathop{\mathrm{Sp}}\nolimits_{2m}(q)$ is the subgroup of $\mathop{\mathrm{GL}}\nolimits_{2m}(q)$ consisting of all isometries of a non-singular alternating bilinear form on the vector space $\mathbb{F}_{q}^{2m}$. It turns out that $|\mathop{\mathrm{Sp}}\nolimits_{2m}(q)|>q^{m^{2}}$. As $|\mathop{\mathrm{GL}}\nolimits_{2m}(q)|=q^{4m^{2}}$, clearly
@@
Thus, the description of the group $\mathop{\mathrm{Sp}}\nolimits_{2m}(q)$ is easy to construct. Given $2m$ and $q$ in binary, in time polynomial in $m$ and $q$ we can define a non-singular alternating bilinear form $\langle$,$\rangle$ , ⟩ on $\mathbb{F}_{q}^{2m}$. Then, as done for the previous case, we can enumerate the elements of $\mathop{\mathrm{GL}}\nolimits_{2m}(q)$, one by one, and put each into $\mathop{\mathrm{Sp}}\nolimits_{2m}(q)$ if it is an isometry for $\langle$,$\rangle$ , ⟩. Now, the simple group $\mathop{\mathrm{PSp}}\nolimits_{2m}(q)$ by quotienting out the subgroup of scalar matrices (which has size at most $2$). Clearly, Lemma 8.2 can be applied to obtain the desired description for $\mathop{\mathrm{PSp}}\nolimits_{2m}(q)$ along with the algorithm for obtaining its multiplication table.
For the next case $\mathop{\mathrm{PSU}}\nolimits_{n}(q)$, we start with the group $\mathop{\mathrm{GU}}\nolimits_{n}(q)$ which is the subgroup of $\mathop{\mathrm{GL}}\nolimits_{n}(q^{2})$ that preserves the sesquilinear form $\langle x,y\rangle=\sum_{i=1}^{n}x_{i}y_{i}^{q}$. From there on the process of obtaining $\mathop{\mathrm{SU}}\nolimits_{n}(q)$ and then $\mathop{\mathrm{PSU}}\nolimits_{n}(q)$ is exactly like in the case of $\mathop{\mathrm{PSL}}\nolimits_{n}(q)$. In order to obtain the desired description for $\mathop{\mathrm{PSU}}\nolimits_{n}(q)$ it suffices to generate $\mathop{\mathrm{GU}}\nolimits_{n}(q)$ in time polynomial in $|\mathop{\mathrm{PSU}}\nolimits_{n}(q)|$ given $n$ and $q$. Clearly, $\mathop{\mathrm{GU}}\nolimits_{n}(q)$ can be generated in time polynomial in $|\mathop{\mathrm{GL}}\nolimits_{n}(q^{2})|$ by looking all these matrices. However, $|\mathop{\mathrm{GL}}\nolimits_{n}(q^{2})|<|\mathop{\mathrm{GU}}\nolimits_{n}(q)|^{3}$. Thus, the brute-force algorithm for generating all of $\mathop{\mathrm{GL}}\nolimits_{n}(q^{2})$ by looking through all $n\times n$ invertible matrices over $\mathbb{F}_{q}$ and picking those which preserve the sesquilinear form will again work. Finally, the group $\mathop{\mathrm{PSU}}\nolimits_{n}(q)$ is only smaller than $\mathop{\mathrm{GU}}\nolimits_{n}(q)$ by a small power of $q$.
@@ 8.1.2 Exceptional Lie type
We will closely follow Wilson’s exposition [79, Chapter 4]. These finite simple groups come in ten families:
@@
    $\mathop{\mathrm{Sz}}\nolimits(2^{2n+1})$:   Let $q=2^{2n+1}$. The Suzuki group $Sz(q)$ (see [79, Chapter 4.2]) is a subgroup of the symplectic group $\mathop{\mathrm{Sp}}\nolimits_{4}(2^{2n+1})$. It arises as the centralizer of a particular automorphism of order $2$ in $\mathop{\mathrm{Sp}}\nolimits_{4}(2^{2n+1})$. As $|\mathop{\mathrm{Sz}}\nolimits(q)|=(q^{2}+1)q^{2}(q-1)$ $|\mathop{\mathrm{GL}}\nolimits_{4}(q)|\leq|\mathop{\mathrm{Sz}}\nolimits(q)|^{4}$. Thus, we can cycle through all elements of $\mathop{\mathrm{Sp}}\nolimits_{4}(q)$ looking for the elements of this centralizer. That automorphism of order $2$ in $\mathop{\mathrm{Sp}}\nolimits_{4}(2^{2n+1})$ is also easily described (in size independent of $q$).
@@
    $G_{2}(q)$:  It turns out that $G_{2}(q)$ (see [79, Chapter 4.3]) arises as the automorphism group of the octonion algebra over $\mathbb{F}_{q}$. The entire octonion algebra can be listed in time $O(q^{8})$. As the automorphisms permute fix the identity in the octonion algebra, $G_{2}(q)$ is essentially a subgroup of $\mathop{\mathrm{GO}}\nolimits_{7}(q)$. We can enumerate $\mathop{\mathrm{GO}}\nolimits_{7}(q)$ and pick out the automorphisms which will give us $G_{2}(q)$. As $|G_{2}(q)|$ is $q^{6}(q^{6}-1)(q^{2}-1)$, this will be polynomial time in the group size. We can define the $\mathop{\mathrm{desc}}(G_{2}(q))$ of size $O(\log q)$. It is essentially just the encoding of $q$.
@@
    $R_{1}(3^{2n+1})={}^{2}G_{2}(3^{2n+1})$: This can be handled quite analogously to $G_{2}(q)$ [79, Section 4.5].
@@
    ${}^{3}{D}_{4}(q)$:  This group (see [79, Chapter 4.6]) arises as centralizer of a certain order $3$ automorphism of $\mathop{\mathrm{P}\Omega}\nolimits^{+}_{8}(q^{3})$. Because of the special nature of this automorphism (product of an $\mathbb{F}_{q}$ automorphism and the triality automorphism), it has a short description of size $O(\log q)$. As $\mathop{\mathrm{P}\Omega}\nolimits^{+}_{8}(q^{3})$ can be enumerated in time polynomial in $q$, we can pick out the centralizer elements in time polynomial in $q$.
@@
    $F_{4}(q)$:  when $q$ is not a power of $2$ or $3$, $F_{4}(q)$ arises as automorphisms of the finite (non-associative) Albert algebra which is $27$-dimensional [79, Chapter 4.8]. Thus, $F_{4}(q)$ can be picked out from $GL_{27}(q)$ by enumerating it and checking each element for being an automorphism. The finite Albert algebra can be defined in $O(\log q)$ size by defining the multiplication of its basis elements.
@@
    $R_{2}(2^{2n+1})={}^{2}F_{4}(2^{2n+1})$:  This case is quite similar to the previous one [79, Section 4.9].
@@
    $E_{6}(q)$:  This group can be defined as automorphisms of a certain trilinear form over the Albert algebra [79, Chapter 4.10]. After quotienting out the scalar subgroup from this automorphism group we will obtain $E_{6}(q)$. The same algorithm of enumerating the bigger group and selecting the automorphisms and then quotienting out the scalars will work.
@@
    $E_{7}(q)$ and $E_{8}(q)$:  Both these arise as automorphisms of a finite Lie algebra [79, Chapter 4.12]. For $E_{8}(q)$ we have to consider a $248$ dimensional Lie algebra. Thus we can look for $E_{8}(q)$ inside $GL_{248}(q)$ by enumerating all its elements and picking out the automorphisms. All we need is that from a $O(\log q)$ size description we should be able to obtain the Lie algebra. For that we only need to express the multiplication of basis elements as an $\mathbb{F}_{q}$ linear combination of the basis (which is of size $248$). $E_{7}(q)$ is similarly dealt with as automorphisms of a smaller dimensional Lie algebra.$\qed$
@@ 8.2 Short descriptions for finite groups
There is a polynomial-time algorithm that takes a group $G$ by its multiplication table as input and outputs a cube-generating sequence of length at most $2\log|G|$ for $G$.
Proof 19
@@
$C_{i}=\{g_{1}^{\epsilon_{1}}g_{2}^{\epsilon_{2}}\cdots g_{i}^{\epsilon_{i}}\mid\epsilon_{j}\in\{0,1\}\hbox{ for }1\leq j\leq i\}.$
For any $g\in G$ let $C_{i}g=\{xg\mid x\in C_{i}\}$. As $|C_{i}|\leq 2^{i}$ and each element in $C_{i}$ can be computed with $i-1$ group multiplications, the set $C_{i}$ can be enumerated with at most $(i-1)2^{i}$ many group multiplications. We can enumerate $C_{i}g$ with $2^{i}$ additional group multiplications.
Now, if $C_{i}g\cap C_{i}\neq\emptyset$ for $g\in G$ then $g$ is in the cube generated by the sequence $(g_{i}^{-1},g_{i-1}^{-1},\ldots,g_{1}^{-1},g_{1},\ldots,g_{i})$. Thus, if for all $g\in G$ $C_{i}g\cap C_{i}\neq\emptyset$ then we have a cube generating sequence of length $2i$. Otherwise, there is a $g_{i+1}\in G$ such that
@@
and the cube $C_{i+1}$ generated by $(g_{1},g_{2},\ldots,g_{i+1})$ is twice the cardinality of $C_{i}$ by a simple counting argument. Clearly, we can find $g_{i+1}$ in time polynomial in $|G|$ by searching through $G$.
As $|C_{i+1}|=2\cdot|C_{i}|$, it follows that this procedure will eventually find a cube generating sequence $C_{k},k\leq 2\log|G|$ with at most polynomial in $|G|$ group multiplications and hence in time polynomial in $|G|$. $\qed$
@@
where $w(x,y)$ is a word over the cube generating sequence $Y_{i-1}$. Thus, $w(x,y)$ is a word of length at most $2\log|N_{i-1}|$, expressed using the c.g.s. $Y_{i-1}$. Let $R_{i}$ consist of the set of all such equations defined by Equation (8) for $y\in Y_{i-1}$ and $x\in X_{i}$. Notice that $R_{i}$ has $|Y_{i-1}|\cdot|X_{i}|$ equations in it and each word $w(x,y)$ over $Y_{i-1}$ is defined by the bit vector of exponents in the cube product over $Y_{i-1}$. Thus, the size of $R_{i}$ is
@@
Using the above claim repeatedly we can express products of the form $xa$ for $x\in C,a\in N_{i-1}$ as $a^{\prime}x$ for some $a^{\prime}\in N_{i-1}$. Finally, we can express the product of any two elements $a_{1}b_{1},a_{2}b_{2}\in N_{i}$, where $a_{1},a_{2}\in N_{i-1}$ and $b_{1},b_{2}\in C$ in the form $a_{3}b_{3},a_{3}\in N_{i-1},b_{3}\in C$. The algorithm runs in time polynomial in $|N_{i-1}|$ and $|X_{i}|$ and hence in time polynomial in $|N_{i}|$.
@@
    More precisely, for each $g\in G$ we compute its normal closure by starting with the cyclic subgroup $N=\langle g\rangle$ and repeatedly replacing it with the (larger) subgroup generated by the set $\{xyx^{-1}\mid x\in G,y\in N\}$ until $N$ does not grow any further. By construction, the final $N$ is a normal subgroup of $G$ and its computation takes time polynomial in $|G|$. If $G$ is not simple, then for some $g\in G$ this construction will yield a non-trivial normal subgroup $N$. We can now recursively find composition series for $N$ and $G/N$ (we have multiplication tables for both) and put them together to obtain a composition series for $G$ in time polynomial in $|G|$.
@@
Our algorithm described in Theorem 8.11 is broadly based on a result of McIver and Neumann [58] that bounds the number of $n$ element groups by $2^{O(\log^{3}n)}$. There is related work by Babai et al. [5] which shows that all simple groups of $n$ elements (except one family called the Ree groups) have generator-relator presentations of size $O(\log^{2}n)$.
@@ 9 Isoclinism of groups and commuting graphs
Motivated by the polynomial-time algorithm for recognizing commuting graphs of extraspecial groups, in this section we briefly discuss isoclinism of groups and its connected to commuting graphs.
$[\theta(g_{1}Z(G)),\theta(g_{2}Z(G))]=\phi([g_{1}Z(G),g_{2}Z(G)]).$
Theorem 9.1
Isoclinic groups of the same order have isomorphic commuting graphs.
Remark 8
It follows from Lemma 7.2 that the extraspecial groups $G_{1}$ and $G_{2}$, defined in Section 7, are isoclinic. Indeed, the recognition algorithm of Section 7 implicitly uses this property.
Is there a converse to Theorem 9.1? That is, if two groups have isomorphic commuting graphs, must they be isoclinic? This holds for various classes of groups, such as abelian groups, non-abelian simple groups, and extraspecial groups (as we have seen).
However, it is not true in general; we recycle an example taken from [23] to show this. Let $G$ be the group of order $64$ which is SmallGroup(64,182) in the SmallGroups library in GAP [37]. A Schur cover of $G$ is a group $H$ of maximal order subject to having a subgroup $Z\leq H^{\prime}\cap Z(H)$ such that $H/Z\cong G$. The Schur multiplier of $G$ has order $2$, so any Schur cover of $G$ has order $128$. Moreover, the Bogomolov multiplier (see [23] for context and other references) of $G$ is equal to the Schur multiplier, which implies that any Schur cover $H$ is commutation-preserving: that is, two elements $a,b\in H$ commute if and only if their projections $Za,Zb\in G$ commute. This implies that the commuting graph of a Schur cover is obtained from the commuting graph of $G$ by replacing each vertex with a clique of size $2$, with all edges between cliques corresponding to adjacent vertices. This procedure also describes the commuting graph of $G\times C_{2}$. On the other hand, it is easy to verify computationally that the derived groups of $H$ and $G\times C_{2}$ are not isomorphic, so these groups cannot be isoclinic.
@@
This is interesting because, as observed, extraspecial groups are nilpotent of class two $2$ and are determined up to isoclinism by their commuting graphs.